You find that Brook Apartment is financed with 30% debt, and the debt investors require a return of 12.5% and equity investors require a 16% return. If the tax rate is 21%, what is the WACC for Brook Apartment?

14.8%

11.8%

14.2%

13.3%
Answer #1

Weight of Debt = D/V = 30%

Weight of Equity = E/V = 70%

Cost of Debt = kd = 12.5%

Cost of Equity = ke = 16%

Tax rate = t = 21%

WACC is calculated using the formula:

WACC = (D/V)*kd*(1-t) + (E/V)*ke = 30%*12.5%*(1 - 21%) + 70%*16% = 2.9625%+11.2% = 14.1625% ~ 14.2% (Rounded to two decimals)

WACC = 14.2%

Answer -> 14.2%